# Enhancing Generalization in Vision–Language–Action Models by Preserving Pretrained Representations

### v1

@@ Abstract
Vision–language–action (VLA) models finetuned from vision–language models (VLMs) hold the promise of leveraging rich pretrained representations to build generalist robots across diverse tasks and environments. However, direct fine-tuning on robot data often disrupts these representations and limits generalization. We present a framework that better preserves pretrained features while adapting them for robot manipulation. Our approach introduces three components: (i) a dual-encoder design with one frozen vision encoder to retain pretrained features and another trainable for task adaptation, (ii) a string-based action tokenizer that casts continuous actions into character sequences aligned with the model’s pretraining domain, and (iii) a co-training strategy that combines robot demonstrations with vision–language datasets emphasizing spatial reasoning and affordances. Evaluations in simulation and on real robots show that our method improves robustness to visual perturbations, generalization to novel instructions and environments, and overall task success compared to baselines.
@@ II Related Work
Figure 2: Method overview. Our approach improves VLA generalization through three designs: (a) Co-training: Jointly training on robotic and vision-language datasets that emphasize spatial affordance and reasoning helps preserve pretrained representations for generalization. (b) Partially-frozen visual encoders: One encoder is frozen to retain robust pretrained features from VLMs, whereas the other is trained to specialize for the robot tasks. (c) String tokenizer: Robotic actions are expressed as digit-based strings to maximally reuse pretrained language representations and to unify prediction targets across non-robotic and robotic task domains for co-training.
@@ IV-B Analysis of Representation Generalization and Robustness
Table IV: Language robustness evaluation. Evaluating policies under original and paraphrased instructions. “w/ Inst. Aug.” indicates training with instruction augmentations.
@@
Language robustness. Table IV compares three settings: (1) baseline VLA models, (2) baselines trained with paraphrased-instruction augmentation, and (3) our models. Unlike (2), which sees all paraphrased evaluation instructions during training, both (1) and (3) see none. Despite receiving no language augmentations, our models substantially outperform baselines trained with augmentation. This shows that our method achieves stronger instruction generalization by leveraging pretrained representations through our string-based action tokenizer, effectively inheriting the language understanding and spatial reasoning capabilities of upstream VLMs.

### v2

@@ Abstract
Vision–language–action (VLA) models finetuned from vision–language models (VLMs) hold the promise of leveraging rich pretrained representations to build generalist robots across diverse tasks and environments. However, direct fine-tuning on robot data often disrupts these representations and limits generalization. We present a framework that better preserves pretrained features while adapting them for robot manipulation. Our approach introduces three components: (i) a dual-encoder design with one frozen vision encoder to retain pretrained features and another trainable for task adaptation, (ii) a string-based action tokenizer that casts continuous actions into character sequences aligned with the model’s pretraining domain, and (iii) a co-training strategy that combines robot demonstrations with vision–language datasets emphasizing spatial reasoning and affordances. Evaluations in simulation and on real robot show that our method improves robustness to visual perturbations, generalization to novel instructions and environments, and overall task success compared to baselines.
@@ II Related Work
Figure 2: Method overview. Our approach improves VLA generalization through three designs: (a) Co-training: Jointly training on robotic and vision-language datasets that emphasize spatial affordance and reasoning helps preserve pretrained representations for generalization. (b) Partially-frozen visual encoders: One encoder is frozen to retain robust pretrained features from VLMs, whereas the other keeps the full flexibility to specialize for the robot tasks. (c) String tokenizer: Robotic actions are expressed as digit-based strings to maximally reuse pretrained language representations and to unify prediction targets across non-robotic and robotic task domains for co-training.
@@ IV-B Analysis of Representation Generalization and Robustness
Table IV: Language robustness evaluation. Evaluating policies under original and paraphrased instructions. “w/ Inst. Aug.” indicates training with instruction augmentations. We also report OpenVLA+SC in this table to show that OpenVLA+ has better robustness, despite performing very similarly for in-distribution evaluation in Tab. II.
@@
Language robustness. Table IV compares three settings: (1) baseline VLA models, (2) baselines trained with paraphrased-instruction augmentation, and (3) our models. Unlike (2), where the paraphrased instructions used for evaluation are disjoint from those seen during training, both (1) and (3) have not seen any paraphrased instructions. The same set of paraphrased instructions is used consistently across all evaluations. Despite receiving no language augmentations, our models substantially outperform baselines trained with augmentation. This shows that our method achieves stronger instruction generalization by leveraging pretrained representations through our string-based action tokenizer, effectively inheriting the language understanding and spatial reasoning capabilities of upstream VLMs.